OpenStudy (anonymous):

What are the solutions of 2x2 - 6x = -8?

OpenStudy (anonymous):

c'mon

OpenStudy (anonymous):

I'm awful i know i know

OpenStudy (anonymous):

divide by 2 because otherwise it will be more annoying, get \[x^2+3x+4=0\] then the damned formula again

OpenStudy (anonymous):

\[x^2-3x+4=0\] is more like it

OpenStudy (anonymous):

yeah you gotta learn how to find \(b^2-4ac\) without lunching it is something any adult can do

OpenStudy (anonymous):

\[a=1,b=-3,c=4\] let me know what you get for \[\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

hint the number inside the radical (the discriminant as we say in the math biz) is negative, so your solutions are complex, i.e. they will have an \(i\) in them

OpenStudy (anonymous):

well i have 3 +/- sort -7/2

OpenStudy (anonymous):

im not sure how to get i or what to do about all of that... @satellite73

OpenStudy (anonymous):

k that is good

OpenStudy (anonymous):

in fact it is completely right we are just going to make it look different \[\sqrt{-7}=\sqrt{7\times (-1)}=\sqrt7\sqrt{-1}=\sqrt7 i\]

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

it's always the same question.

OpenStudy (anonymous):

note that you do all those steps in your head i just wrote them out to show what i was thinking not what you need to write

OpenStudy (anonymous):

i know, but ya know.. girls gotta get help haha @RaphaelFilgueiras

OpenStudy (anonymous):

final answer \[\frac{3\pm\sqrt7 i}{2}\]

OpenStudy (anonymous):

note that the \(i\) is OUTSIDE the radical if you write \[\large \sqrt{7i}\] your teacher will think you are a moron

OpenStudy (anonymous):

my teacher is a moron herself so i don't think it will be a problem (: but ill be sure its outside of the radical haha @satellite73

OpenStudy (anonymous):

no doubt who takes algebra in august?